OpenStudy (anonymous):

Why does ln(1/2) = -ln(2) please explain mechanic to get to this solution

OpenStudy (mertsj):

1/2 = 2^-1

OpenStudy (anonymous):

several reasons. one is that \[\log(b^n)=n\log(b) \]and therefore \[\ln(2^{-1})=-1\ln(2)\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

why does ln(1/4) not equal -ln4 ??

OpenStudy (mertsj):

It does, Frank. Check it out on your calculator

OpenStudy (anonymous):

what mertsj said. it certainly does

OpenStudy (anonymous):

my calculator says -2ln(2)

OpenStudy (anonymous):

any idea as to why, my calculator is a ti 89 titanium